the number of atoms present in 215 g of silver nitrate, AgNO3
1
Expert's answer
2020-04-18T07:07:09-0400

N=(m(AgNO3)/Mr(AgNO3) )*Na=(215/170)*6.02*10^23=7.613*10^23